# captionUnknown document class

### v1

@@ Abstract
Randomized benchmarking (RB) is the most commonly employed protocol for the characterization of unitary operations in quantum circuits due to its reasonable experimental requirements and robustness against state preparation and measurement (SPAM) errors. So far, the protocol has been limited to discrete or fermionic systems, whereas extensions to bosonic systems have been unclear for a long time due to challenges arising from the underlying infinite-dimensional Hilbert spaces. In this work, we close this gap and develop an RB protocol for passive Gaussian transformations, which we call bosonic passive RB. The protocol is based on the recently developed filtered RB framework [helsenGeneralFrameworkRandomized2020, heinrichRandomizedBenchmarkingRandom2023] and is designed to isolate the multitude of exponential decays arising for bosonic systems. We give explicit formulas and a Julia implementation for the necessary post-processing of the experimental data. We also analyze the sampling complexity of bosonic passive RB by deriving analytical expressions for the variance. They show a mild scaling with the number of modes, suggesting that bosonic passive RB is experimentally feasible for a moderate number of modes. We focus on experimental settings involving Fock states and particle number resolving measurements, but also discuss Gaussian settings, deriving first results for heterodyne measurements.
@@ I Introduction
Bosonic quantum systems play a major role in the design of quantum computing platforms. Most prominently, this includes photonic quantum computing as a popular proposal for scalable quantum computers [knillSchemeEfficientQuantum2001, koashiProbabilisticManipulationEntangled2001, kokLinearOpticalQuantum2007, rudolph2016i, tanResurgenceLinearOptics2018, alexanderManufacturablePlatformPhotonic2024]. The biggest advantages of this model rely on the implementation of particle sources, detectors, and linear optical circuits on the same integrated chips [politiSilicaonSiliconWaveguideQuantum2008, sprengersWaveguideSuperconductingSinglephoton2011, silverstoneOnchipQuantumInterference2014, meanyLaserWrittenCircuits2015], and access to mixed schemes for quantum error correction [bourassaBlueprintScalablePhotonic2021]. Bosonic systems also offer interesting non-universal models of computation to test quantum supremacy, such as boson sampling [aaronsonComputationalComplexityLinear2010] and Gaussian boson sampling [hamiltonGaussianBosonSampling2017, chakhmakhchyanBosonSamplingGaussian2017], which have been implemented successfully in real experiments recently [zhongQuantumComputationalAdvantage2020, zhongPhaseProgrammableGaussianBoson2021, madsenQuantumComputationalAdvantage2022].
Most challenges in the characterization of bosonic systems can be attributed to the particularities of the infinite-dimensional Hilbert space [lvovskyContinuousvariableOpticalQuantum2009, sharmaCharacterizingPerformanceContinuousvariable2020]. For instance, protocols that rely on scrambling techniques via unitary designs are challenging to adapt, since Gaussian unitaries only form a unitary $1$-design [blume-kohoutCuriousNonexistenceGaussian2011, zhuangScramblingComplexityPhase2019], and unitary $2$-designs cannot exist for CV systems, unless rigged Hilbert spaces are considered [iosueContinuousvariableQuantumState2022]. Such issues constrain characterization protocols to very specific settings, where the implemented unitary operators (which may include non-Gaussian single-mode unitaries) are only benchmarked w.r.t. a restricted set of input states [fariasCertificationContinuousvariableGates2021].
@@
In this work, we introduce a RB protocol for bosonic systems, bosonic passive RB (or passive RB for simplicity), for the benchmarking of passive (Gaussian) transformations. To this end, we build on previous work on filtered RB [heinrichRandomizedBenchmarkingRandom2023], and make the common assumption of gate-dependent, stationary, and Markovian noise. To circumvent problems arising from infinite dimensions, we consider experiments where the input state is a number state and PNR (PNR) measurements are performed, resembling boson sampling experiments [aaronsonComputationalComplexityLinear2010]. For practical reasons, we propose to use collision-free states with at most one particle per mode. We then show that the passive RB succesfully benchmarks the average quality of passive transformations on a fixed number of particles, thereby separating the effect of particle number-preserving noise from particle loss using different post-processing of the same experimental data. In general, we prove that there are as many decay rates as we have particles in the experiments – however it is sufficient to only estimate a constant number of them in practice. By representation-theoretic means, we derive explicit and concise formulas for the functions used in the post-processing and provide a Julia implementation. We simulate the passive RB protocol for a small number of modes, demonstrating that it works as intended. Finally, we analyze the sampling complexity of our protocol by computing the relevant variances. We evaluate the obtained expressions numerically and find a very mild scaling with the number of modes $m$ which may be logarithmic – a definite conclusion is however prevented by the small number of data points.
The effectively finite-dimensional setting is introduced for a good reason: Although arguably a simpler experimental setting, Gaussian input states and measurements result in infinitely many relevant irreps and thus decay rates, rendering a meaningful analysis impossible. As a first step, we also consider scenarios in which the input state is still a number state, but each mode is measured with a (balanced) heterodyne detector. In this cases, passive RB is similar to Gaussian boson sampling experiments [hamiltonGaussianBosonSampling2017, chakhmakhchyanBosonSamplingGaussian2017].

### v2

@@ Abstract
Randomized benchmarking (RB) is the most commonly employed protocol for the characterization of unitary operations in quantum circuits due to its reasonable experimental requirements and robustness against state preparation and measurement (SPAM) errors. So far, the protocol has been limited to discrete or fermionic systems, whereas extensions to bosonic systems have been unclear for a long time due to challenges arising from the underlying bosonic Hilbert space. In this work, we close the gap for bosonic systems and develop an RB protocol to benchmark passive Gaussian transformations on any particle number subspace, which we call passive bosonic RB. The protocol is built on top of the recently developed filtered RB framework [helsenGeneralFrameworkRandomized2020, heinrichRandomizedBenchmarkingRandom2023] and is designed to isolate the multitude of exponential decays arising for passive bosonic transformations. We give explicit formulas and a Julia implementation for the necessary post-processing of the experimental data. We also analyze the sampling complexity of passive bosonic RB by deriving analytical expressions for the variance. They show a mild scaling with the number of modes, suggesting that passive bosonic RB is experimentally feasible for a moderate number of modes. We focus on experimental settings involving Fock states and particle number resolving measurements, but also discuss Gaussian settings, deriving first results for heterodyne measurements.
@@ I Introduction
Bosonic quantum systems play a major role in the design of quantum computing platforms. Most prominently, this includes photonic quantum computing as a popular proposal for scalable quantum computers [knillSchemeEfficientQuantum2001, koashiProbabilisticManipulationEntangled2001, kokLinearOpticalQuantum2007, rudolph2016i, tanResurgenceLinearOptics2018, alexanderManufacturablePlatformPhotonic2024]. The biggest advantages of this model rely on the implementation of particle sources, detectors, and linear optical circuits on the same integrated chips [politiSilicaonSiliconWaveguideQuantum2008, sprengersWaveguideSuperconductingSinglephoton2011, silverstoneOnchipQuantumInterference2014, meanyLaserWrittenCircuits2015], and access to mixed schemes for quantum error correction [bourassaBlueprintScalablePhotonic2021]. Bosonic systems also offer interesting non-universal models of computation to test quantum supremacy, such as (Gaussian) boson sampling [aaronsonComputationalComplexityLinear2010, hamiltonGaussianBosonSampling2017, chakhmakhchyanBosonSamplingGaussian2017], which have been implemented successfully in real experiments recently [zhongQuantumComputationalAdvantage2020, zhongPhaseProgrammableGaussianBoson2021, madsenQuantumComputationalAdvantage2022].
Most challenges in the characterization of bosonic systems can be attributed to the particularities of the bosonic Hilbert space [lvovskyContinuousvariableOpticalQuantum2009, sharmaCharacterizingPerformanceContinuousvariable2020]. For instance, protocols that rely on scrambling techniques via unitary designs are challenging to adapt, since Gaussian unitaries only form a unitary $1$-design [blume-kohoutCuriousNonexistenceGaussian2011, zhuangScramblingComplexityPhase2019], and unitary $2$-designs cannot exist for CV systems, unless rigged Hilbert spaces are considered [iosueContinuousvariableQuantumState2022]. Such issues constrain characterization protocols to very specific settings, where the implemented unitary operators (which may include non-Gaussian single-mode unitaries) are only benchmarked w.r.t. a restricted set of input states [fariasCertificationContinuousvariableGates2021].
@@
In this work, we introduce a RB protocol for the benchmarking of passive bosonic (Gaussian) transformations, called passive bosonic RB, or short passive RB. The passive RB protocol can be understood as a bosonic incarnation of the filtered RB framework [heinrichRandomizedBenchmarkingRandom2023], and we build on its results to give a comprehensive analysis of passive RB, including its sampling complexity, under the assumption of gate-dependent, stationary, and Markovian noise.
In contrast to the discrete variable systems explicitly considered in Ref. [heinrichRandomizedBenchmarkingRandom2023], reasonable guarantees for a filtered RB protocol in bosonic systems are harder to obtain. This is because the action of passive transformation on the full bosonic Hilbert space is considerably more complicated than, e.g., the action of the qubit Clifford group, and thus a well-behaved protocol requires a careful identification of relevant and feasible experimental settings.
@@
In the passive RB protocol, we navigate these complications by considering experiments where the input state is a number state and PNR (PNR) measurements are performed, resembling boson sampling experiments [aaronsonComputationalComplexityLinear2010]. For practical reasons, we propose to use collision-free states with at most one particle per mode. We then show that the passive RB successfully benchmarks the average quality of passive transformations on a fixed number of particles, thereby separating the effect of particle number-preserving noise from particle loss using different post-processing of the same experimental data. In general, we prove that there are as many decay rates as we have particles in the experiments – however it is sufficient to only estimate a constant number of them in practice. By representation-theoretic means, we derive explicit and concise formulas for the functions used in the post-processing and provide a Julia implementation. As the protocol involves non-Gaussian elements, the classical post-processing requires the computation of matrix permanents, and it is therefore generally inefficient [aaronsonComputationalComplexityLinear2010]. Specifically, we expect the post-processing to be feasible for $\approx 30$ particles and modes, thus allowing the protocol to target an experimentally interesting regime. The inefficiency of the post-processing should not be unexpected, as a similar behaviour can be found in linear cross-entropy benchmarking [liu_benchmarking_2022, heinrichRandomizedBenchmarkingRandom2023]. There, the post-processing involves the simulation of random circuit sampling, which is also known to be computationally hard [boixoCharacterizingQuantumSupremacy2018, aruteQuantumSupremacyUsing2019, liuRedefiningQuantumSupremacy2021, hangleiter_computational_2023]. We simulate the passive RB protocol for a small number of modes, demonstrating that it works as intended. Finally, we analyze the sampling complexity of our protocol by computing the relevant variances. We evaluate the obtained expressions numerically and find a very mild scaling with the number of modes $m$ which seems be logarithmic.
The effectively finite-dimensional setting is introduced for a good reason: Probing the full bosonic Hilbert space involves infinitely many relevant irreps and thus also decay rates which, a priori, render a meaningful analysis impossible. Moreover, such irreps would also appear with non-trivial multiplicities, making the protocol numerically challenging from a practical point of view, as it would require fitting matrix exponential decays [helsenGeneralFrameworkRandomized2020, heinrichRandomizedBenchmarkingRandom2023]. As a first variation of the passive RB setup, we thus consider scenarios in which the input state is still a number state, but the measurement involves (balanced) heterodyne detectors.